cerna.gif (1167 bytes)

LBOC Meetings 2015

all meetings on INDICO

lhc_logo.gif (4175 bytes)


Topics presented

Presence sheet
20 January 2015
  • Luminosity, beam loss, and emittance growth revisited again (M. Lamont)
  • Status of optics for 2015 (R. De Maria)
  • Status of 2015 optics in LSA (J. Wenninger)
3 February 2015
  • Building a Luminosity model for the LHC and HL-LHC (F. Antoniou)
  • Secondary beam losses in Pb-Pb operation in 2011 and perspective for 2015 (M. Schaumann)
  • Considerations on ALICE BCM dump thresholds for RUN 2 (A. Di Mauro)
17 February 2015
  • Emittance Measurement Preparation during LHC Commissioning (M. Kuhn)
  • Status of optics and cycles in LSA (J. Wenninger)
24 February 2015
  • Planning of the sector test (R. Alemany)
  • Status of dry runs & machine checkout (R. Alemany and J. Wenninger)
10 March 2015
  • Brief summary of the sector test (V. Kain)
  • IP1, IP2, IP5 and IP8 crossing angle recommendations (T. Pieloni)
  • Beam-based Modelling (P. Skowronski)
17 March 2015
  • Machine Check-out status and plan (J. Wenninger)
  • Update on faster ramp-down (M. Solfaroli Camillocci)
  • Status of Beam Instrumentation (E. Bravin)
  • Status of collimator commissioning (B. Salvachua Ferrando)
14 April 2015
  • Beam dumps at 6.5 TeV (J. Wenninger)
  • LHC configuration (J. Wenninger)
  • Commissioning status of RF (P. Baudrenghien)
  • First optics results (T. Persson)
21 April 2015
  • Status of emittance measurements (E. Bravin)
  • Status of wire-scanner and k-modulation (M. Kuhn)
5 May 2015
  • Beam transfer function measurements (T. Pieloni)
  • Proposal for margin allocations for 2015 startup at 80cm (S. Redaelli)
19 May 2015
  • Status of Collimation at 450 GeV (G. Valentino)
  • Investigation of Unidentified Lying Object in C15R8 (D. Mirarchi)
26 May 2015
  • Monitoring with BSRT (G. Trad)
  • MPS considerations and intensity ramp up (D. Wollmann)
  • Monitoring of heat load (J. Wenninger)
  • Scrubbing run planning (G. Iadarola)
2 June 2015
  • Injection considerations - MKI TDI (C. Bracco)
  • Vacuum considerations (G. Bregliozzi, C. Yin Vallgren)
  • Cryo considerations (L. Tavian)
  • Stable phase shift measurement (E. Shaposhnikova)
16 June 2015
  • Chromaticity and b3: decay, snapback & flat top (M. Solfaroli Camillocci)
  • Chromaticity Reprocessing: Fill 3806 (L. Carver)
  • Overview of aperture measurements (P. Hermes)
23 June 2015
  • Doublet beams in the LHC: BPM system (T. Lefevre)
  • Doublet beams in the LHC: RF (H. Tinko)
  • Doublet beam injection (W. Bartmann)
30 June 2015
  • ADT damping times (W. Hofle)
  • Summary of the 50 ns scrubbing run (G. Iadarola)
07 July 2015
  • Tune decay and snapback (M. Schaumann)
  • Fixed Displays for the follow-up of the beam-induced RF heating (B. Salvant)
28 July 2015
  • LHC emittance measurements and preservation (M. Kuhn)
  • Observations from OP extended luminosity scans (M. Hostettler)
  • A first look at the 2015 fault statistics (A. Apollonio)
18 August 2015
  • Overview on LHC Scrubbing Run 2 (H. Bartosik)
  • Chromaticity correction without RCS.A78B2 (M. Solfaroli)
1 September 2015
  • Summary of instability observations at injection (K. Li)
  • Summary of instability observations at flat top and beginning of squeeze, i.e. without beam-beam (L.R. Carver)
  • Summary of instability observations at end of squeeze and after, i.e. with beam-beam (T. Pieloni )
8 September 2015
  • Cryo availability for operation and scubbing (G. Ferlin)
  • Proposal for an ion squeeze (M. Schaumann)
  • Status and prospects for the RF phase measurements (J. Esteban Muller)
22 September 2015
  • First experience with the Q1 DOROS BPMs (A. Gorzawski)
  • Observation of ground motion in the LHC at the triplet (M. Fitterer)
  • Status of the instability trigger (T. Levens)
6 October 2015
  • Update on beam size measurements from OP scans (M. Hostettler)
  • Observations of instabilities during OP scans (T. Pieloni)
  • Margins to increase ADT gain at injection (W. Hofle)
  • Exploration of Q' and octupole settings at injection (L. Carver)
13 October 2015
  • On possible sources of luminosity differences between ATLAS and CMS (R. Tomas)
  • Crossing angles from K-modulation (J. Wenninger)
  • Longitudinal instability due to bunch shortening in long fills (J. Esteban Muller)
27 October 2015
  • Update on crossing angles in IR1 and IR5 (J. Wenninger)
  • Configuration for the 2.51 TeV pp reference run
  • Effect of the e-cloud on the tune footprint at 450 GeV (A. Romano)
  • Beam induced RF heating in LHC in 2015 (B. Salvant)
24 November 2015
  • Updated results on triplet k-modulation (M. Kuhn)
  • Beta* correction strategies (T. Persson)
  • First observations with ADT obsbox (L. Carver)
8 December 2015
  • LBS summary (R.Alemany)
  • ADT parameter extraction: an alternative approach (G.Kotzian)